### v2

# UltraFusion: Ultra High Dynamic Imaging using Exposure Fusion

Abstract
Capturing high dynamic range (HDR) scenes is one of the most important issues in camera design. Majority of cameras use exposure fusion technique, which fuses images captured by different exposure levels, to increase dynamic range. However, this approach can only handle images with limited exposure difference, normally 3-4 stops. When applying to very high dynamic scenes where a large exposure difference is required, this approach often fails due to incorrect alignment or inconsistent lighting between inputs, or tone mapping artifacts. In this work, we propose UltraFusion, the first exposure fusion technique that can merge input with 9 stops differences. The key idea is that we model the exposure fusion as a guided inpainting problem, where the under-exposed image is used as a guidance to fill the missing information of over-exposed highlight in the over-exposed region. Using under-exposed image as a soft guidance, instead of a hard constrain, our model is robust to potential alignment issue or lighting variations. Moreover, utilizing the image prior of the generative model, our model also generates natural tone mapping, even for very high-dynamic range scene. Our approach outperforms HDR-Transformer on latest HDR benchmarks. Moreover, to test its performance in ultra high dynamic range scene, we capture a new real-world exposure fusion benchmark, UltraFusion dataset, with exposure difference up to 9 stops, and experiments show that UltraFusion can generate beautiful and high-quality fusion results under various scenarios. An online demo is provided at project page.
Figure 1: Comparing our guided inpainting HDR imaging method with a state-of-the-art HDR reconstruction [26] and multi-exposure fusion [49] methods. Both scenes are selected from our captured new real-world benchmark. Left: night cityscape with large exposure difference. Right: afternoon street with motion-induced occlusion. Previous methods struggle to handle these scenes. By modeling HDR as inpainting problem, our method can produce visually appealing results without ghosting artifacts in these challenging scenes.
1 Introduction
High dynamic range (HDR) imaging is one of the fundamental problems in the modern camera design. Due to the hardware limitation, camera sensors have a much smaller dynamic range compared with real-world. To increase it, majority of HDR solutions merge multiple images, either same [9] or different exposure levels [15, 50, 56, 57, 35, 3, 25, 41, 44, 17]. Despite all recent advances, majority HDR algorithms can only bring limited increase in dynamic range. For example, HDR+ [9], the first HDR algorithm used by commercial cameras, can only robustly increase the dynamic range by 8 times (3 stops). Therefore, in this work, we study the following question: can we drastically increase the dynamic range of a camera by fusing two images with very large exposure difference, like 9 stops in Fig. 1?
This is a fundamental challenging problem, due to the following three issues. First, to handle dynamic scenes, most of HDR fusion algorithms will first align input frames, which is very challenging when input has large brightness difference. As result, a ghosting issue happens when alignment fails, indicated by zoom-in patches in the right scene of Fig. 1. Second, most of HDR algorithms assume that under-exposed image is simply a darker version of the normal image. However, the appearance of an object may change when exposure levels change, like ship in the left scene of Fig. 1, resulting unnatural fusion result. Third, sometimes the result of fusion is an HDR image, which cannot be directly shown on normal low-dynamic-range display. Therefore, these HDR images will be further compressed through a tone-mapping process. When dynamic range is high, tone-mapping may introduce additional issue. Maintaining a natural contrast and rich details in the final output is challenging, as shown in the zoom-in patches of previous HDR reconstruction method in Fig. 1.
Figure 2: Visual comparison on directly utilizing ControlNet [64] and our UltraFusion. ControlNet struggles to fix a frame for reference, while our method fixes the over-exposed image as the reference and the under-exposed image as the guidance for inpainting, thereby avoiding artifacts.
In this work, we propose a completely different fusion method, UltraFusion, which models it as a guided inpainting problem. In this setup, the user captures two images, one normal exposed image where brighter objects are over-exposed, and another under exposed image, which only captures the very brighter parts of the scene. We use the normal exposed image as a reference, and inpainting the missing information in the highlight. Unlike the traditional inpainting, we use the information from under-exposed frame as guidance, so inpainted highlight is not completely generated, but stays consistent with under-exposed frame.
There are three advantages of this approach when handling large exposure difference. First, it follows the exposure fusion [31] setup. Unlike HDR fusion techniques, which first generate HDR image and then compress to a low-dynamic-range (LDR) image through tone mapping, exposure fusion directly generates LDR output, avoiding cascading errors. Second, the under-exposed image is used as a soft guidance, instead of a hard constraint. Therefore, UltraFusion is robust to alignment error (see the right scene of Fig. 1) and lighting variation (see the left scene of  Fig. 1). Third, image prior of the generative model ensures the natural look of the output image, reducing the potential artifacts.
To train a guided inpainting, a simple solution is to train a ControlNet [64] using two pre-aligned input images. However, such paradigm cannot handle dynamic HDR scenes, as ControlNet may not know which frame to choose as the reference, increasing the difficulty to fuse the information from over and under-exposed images. For instance, as shown in Fig. 2, ControlNet selects the over-exposed image as the reference frame in red boxed region, but chooses the under-exposed image as the reference frame in green boxed region, leading to substantial artifacts in the result. Additionally, as a generative model, it may inevitably generate fake image content, as indicated by green box in Fig. 2.
To address these challenges, the proposed UltraFusion is designed as follows. First, we warp the short-exposed image to the long-exposed one, and mask out the occluded regions. Then, we utilize the diffusion prior to inpaint the long-exposed image guided by partial short-exposed information. To make the network retain more details to generate guidance information, we propose a new decompose-and-fuse control branch, which eliminates the luminance component of the short-exposed image, extracting structure and color information instead, and employs a multi-scale cross-attention to improve the feature fusion with long-exposed image.
To evaluate the effectiveness of our framework, we capture 100 under/over-exposed image pairs, covering daytime, nighttime, indoor, outdoor scenes with local and global motion patterns. Experiments on both latest HDR imaging datasets and our captured benchmark demonstrate that, comparing to existing methods, our UltraFusion is more robust to scenes with large exposure differences and large motion, as shown in Fig. 1.
2 Related work
2.1 HDR imaging
HDR Imaging can be devided into HDR reconstruction and Multi-Exposure Fusion (MEF) typically depends on the domain where fusion occurs [20]. HDR reconstruction methods [15, 50, 56, 57, 35, 3, 25, 41, 44, 52, 17] inverts the camera response function (CRF) to merge exposure brackets in the linear HDR domain [29]. In most cases, tone mapping is necessary to display the reconstructed HDR image properly on standard LDR monitors,. As a cost-effective alternative [13], MEF methods [31, 36, 29, 34, 20, 53, 23, 49, 13, 67, 68] directly fuse images in the LDR domain, sidestepping CRF calibration and sophisticated tone mapping process [31]. Regardless of the type of HDR Imaging methods, they contend with ghosting artifacts caused by camera shake and object movement [40]. Previous methods have attempted explicit or implicit alignment using optical flow [15, 34, 35, 50] or attention mechanism [3, 25, 41, 44, 56, 57, 17]. However, most HDR Imaging methods suffer from unpleasing artifacts when large motion causes occlusion in the complementary region.
Diffusion models. Recently, it has been witnessed with the rapid rise of diffusion models [10, 38] and their successful application in various tasks, including controllable image generation [59, 64], image restoration [32, 22, 48, 24, 62], image editing [30, 4, 39, 14] and image inpainting [27, 5, 51, 63]. In the field of HDR imaging, the application of diffusion models has primarily focused on HDR deghosting [58, 8, 12]. However, since these methods do not levearge the diffusion priors learned from large-scale datasets, their generalization ability is limited by the scale of the HDR dataset. While some recent works [19, 6] have employed diffusion priors, they tend to focus on single image HDR. Without another differently-exposed image as reference, the results generated by these methods lack sufficient reliability. Unlike previous methods, we utilize diffusion priors and use the short-exposed image as a reference to perform reliable inpainting in the highlight regions of the over-exposed image, thereby achieving natural and reliable HDR scene reconstruction. Compared to diffusion-based inpainting methods [27, 5, 51, 63] that perform inpainting solely from scratch, we leverage information from short-exposure images to guide a more accurate inpainting process.
Tone mapping methods. The goal of tone mapping is to convert HDR images to LDR for display on standard screens while enhancing visual detail. Due to the challenge of obtaining ground-truth tone mapping results, unsupervised deep learning approaches have been developed using adversarial [46] and contrastive [2] learning. To address data limitations, Cai et al. [1] manually curated training data by selecting the best results from 13 tone mapping methods for network learning [1, 11]. However, previous tone mapping algorithms, lacking robust image priors and facing data constraints, struggle with visually pleasing results and generalization in extreme high dynamic range scenes. By incorporating diffusion-based image priors, our method achieves aesthetic results even in challenging high dynamic range scenarios (see Fig. 1).
Figure 3: The whole backbone of UltraFusion. Our method is a 2-stage framework, consisting of (a) pre-alignment stage and (b) guided inpainting stage. The first stage pre-aligns the under-exposed image $I_{ue}$ to the over-exposed image $I_{oe}$ and masks the occluded regions. In the subsequent guided inpainting stage, we propose a new decompose-and-fuse control branch to utilize the diffusion priors.
3 Methodology
Given an over-exposed image $I_{oe}$ and an under-exposed image $I_{ue}$, traditional exposure fusion algorithms directly aggregate different frequency band of both images, which are sensitive to misalignment error or lighting variation. Instead, we treat this as an inpainting problem. Specifically, we use the over-exposed image $I_{oe}$ as the base image and inpainting missing information in the highlight region. To ensure inpainted highlights are real, we also use highlights from under-exposed image as guidance.
Based on this idea, we design a 2-stage newtwork shown in Fig. 3, which consists of the pre-alignment stage and the guided inpainting stage. The pre-alignment stage outputs a coarse-aligned version of $I_{ue}$, which is used as the soft guidance in the following guided inpainting stage. Details of each stage is described below.
3.1 Pre-alignment stage
Most of optical flow alignment assumes input have the similar brightness. Therefore, we first adjust the brightness $I_{ue}$ to match the distribution of $I_{oe}$ through intensity mapping function [7]. Then, we adopt RAFT [43], a pre-trained optical flow network, to estimate the bidirectional flow $f_{oe\rightarrow ue}$ and $f_{ue\rightarrow oe}$ and align $I_{ue}$ to $I_{oe}$ using backward warping. However, backward warping will result in ghosting at the occlusion boundary [66], leading to artifacts in the next guided inpainting stage. To solve that, we utilize the forward-and-backward consistency check [55] to estimate occluded regions $\mathcal{M}$ and mask them out in the warped output. Finally, we can obtain a pre-aligned output $I_{ue\rightarrow oe}$ of the first stage:
$I_{ue\rightarrow oe}=(1-\mathcal{M})\cdot\mathcal{W}(I_{ue},f_{oe\rightarrow ue}),$ (1)
where $\mathcal{W}$ denotes the backward warping. Fig. 3 (a) shows the output is a masked and aligned under-exposed image.
3.2 Guided inpainting stage
We build our guided inpainting model based on the Stable Diffusion Model [37] because its powerful generative prior can help to resolve ambiguity during inpainting. Similar to other diffusion-based image enhancement techniques [22], we also inject the following information through an additional control branch, as shown in Fig. 3(b): 1) the image to be inpainted, which is the over-exposed image $I_{oe}$, 2) the additional guidance of highlight, which is the under-exposed image $I_{ue}$, and 3) the diffusion latent at the current diffusion step $z_{t}$, as previous work [22] shows that including diffusion latent as an additional condition can improve image quality. The main diffusion denoising network is a pretrained U-Net. The over-exposed image is first encoded using a pretrained VAE before entering the diffusion module, and the outputs are converted back to the image space using the pretrained decoder.
Different from general diffusion-based image enhancement, we propose a novel decompose-and-fuse control branch to inject two input images and diffusion latent as a control signal, as Fig. 10 (b) shows that naively injecting this information may not be able to guide diffusion to faithfully inpaint missing highlights obtained from the under-exposed image.
Figure 4: The detailed architecture of our proposed decompose-and-fuse control branch.
Decompose-and-fuse control branch. Fig. 4 shows our control branch. We use the over-exposed image $I_{oe}$ as the main control signal and the under-exposed image $I_{ue}$ as the soft guidance. Following ControlNet [64], we copy the encoder and middle blocks of the denoising U-Net as the main extractor, but update their weights during training. A simple soft guidance is to use the encoded under-exposed image latent $y_{ue}$ from the VAE encoder, combined with the over-exposed latent $y_{oe}$. However, the underexposed images are often too dark to be directly used as soft guidance, since the model may entirely ignore this guidance.
Therefore, we decompose the underexposed image into the color and structure information, both of which are robust to brightness changes. Specifically, we use the normalized image as the structure component, similar to SSIM [47], as:
$S_{ue}=(Y_{ue}-\mu(Y_{ue}))/\sigma(Y_{ue}),$ (2)
where $Y_{ue}$ represents the luminance channel of $I_{ue}$ under YUV space, $\mu(\cdot)$ and $\sigma(\cdot)$ denote the mean intensity and standard deviation, respectively. The chroma channels (UV) are used as color information. The extracted structure and color information are further encoded using trained color and structure extractors (gray GE block in Fig. 4).
At last, the extracted features are injected into the main extractor with a multi-scale cross attention, as shown in the bottom part of Fig. 4. The output of each cross attention module is fed into both the next level of the main extractor and the corresponding U-Net block, using zero convolution.
4 Experiment
Figure 5: (a) The data distribution of our benchmark. (b) The user study result on our benchmark.
Table 1: Quantitative comparisons on the static MEFB dataset [65].
[…]
Figure 6: Trade-off curve between MEF-SSIM and MUSIQ on MEFB dataset [65]. Our UltraFusion achieves the best trade-off between image quality and information preservation.
Table 2: Quantitative comparisons on dynamic RealHDRV dataset [40] and our challenging UltraFusion benchmark.
[…]
4.1 Experimental setting
Datasets. We evaluate our method on both static datasets and dynamic datasets. For evaluation, we use the MEFB dataset [65] with 100 static under/over-exposed image pairs and RealHDRV [40], a dynamic HDR deghosting test set containing 50 scenes with varying motion patterns.
UltraFusion benchmark. Existing exposure fusion benchmark cannot fully evaluate fusing in real-world challenging conditions, as they either lack realistic motion [65] or have limited dynamic range [15, 45, 40]. Therefore, we collect a new real-world UltraFusion benchmark, which contains 100 real-captured under/over-exposed image pairs. Compared to previous datasets, our benchmark is more challenging for three reasons: 1) Our benchmark features larger exposure differences between the two input images (up to 9 stops). 2) It includes more realistic motion, with many scenes containing extensive and unintentional foreground movement. 3) Our benchmark is highly diverse, encompassing daytime, nighttime, indoor, and outdoor scenes captured by DSLR Camera (Canon R8) and mobile phones (iPhone12, iPhone13, Redmi K50 Pro and OPPO Reno8 Pro). We summarize the exposure difference distribution and exposure time distribution of our benchmark in Fig. 5 (a). It can be observed that our benchmark covers a wide range of exposure differences and diverse exposure times, which can be used to effectively test the robustness of the HDR methods.
Implementation details. We leverage the generative prior encapsulated in Stable Diffusion V2.1 [38]. The decompose-and-fuse control branch (DFCB) is trained for 140k iterations with batch size $bs=32$ on 8 NVIDIA RTX 4090 GPUs. Adam is adopted as the optimizor and the learning rate is fixed to 0.0001. To adapt HDR reconstruction methods to 2 differently exposed inputs, we re-implement them by following their default settings.
Evaluation metrics. We utilize three widely-used non-reference metric MUSIQ [16], PAQ2PIQ [61] and HyperIQA [42] for quantitative comparison. Moreover, For static dataset [65], as no ground truths are provided, we select the task-specific MEF-SSIM [28] for structure retention evaluation. For dynamic dataset [40] with HDR ground truths, TMQI [60] is used to evaluate the performance of fidelity and naturalness. In addition, we conduct a user study on our Ultrafusion benchmark to perform subjective evaluation.
Figure 7: Visual comparisons of different exposure fusion methods on static MEFB dataset [65].
Figure 8: Visual comparisons on our captured UltraFusion benchmark.
Figure 9: Visual results on dynamic RealHDRV dataset [40].
4.2 Comparisons with HDR imaging methods
We compare our method with state-of-the-art HDR Imaging methods, including HDR reconstruction methods HDR-Transformer [26], SCTNet [45], SAFNet [18], and multi-exposure fusion methods Deepfuse [36], MEF-GAN [54], U2Fusion [53], Defusion [21], MEFLUT [13], HSDS-MEF [49], TC-MoA [68]. As HDR reconstruction methods cannot output an LDR image directly, we use professional software Photomatix [33] to perform tone mapping.
Evaluation on static dataset. We evaluate the fusion performance on the MEFB dataset [65], focusing on large exposure differences. As shown in Fig. 6, our method outperforms other methods on all three non-reference metrics (MUSIQ, PAQ2PIQ, and HyperIQA). Specifically, our proposed UltraFusion achieves 2.06 gain in terms of MUSIQ compared to HSDS-MEF. In terms of MEF-SSIM, as shown in  Fig. 6, our baseline model (ControlNet [64]) achieves better image quality but lacks fidelity, while HSDS-MEF [49] retains more information from inputs at the cost of quality. Our UltraFusion outperforms most algorithms, and achieves similar fidelity scores compared to HSDS-MEF and TC-MoA but with much higher image quality (non-reference metrics), indicating the best trade-off between visual quality and information preservation. The qualitative comparison in Fig. 7 further validates this. In contrast, HDR reconstruction methods (HDR-Transformer and SCTNet) often miss some detail in highlights and MEF methods (HSDS-MEF and TC-MoA) introduce unnatural transition from brighter to dark regions.
Evaluation on dynamic dataset. To further illustrate the robustness of UltraFusion to global and local motion, we use the RealHDRV dataset [40]. We extract the corresponding over-exposed image from the HDR ground truth as input.  Tab. 2 demonstrates that our UltraFusion achieve state-of-the-art performance in terms of all metrics. For dynamic scenes, TMQI metric is a particularly important metric, as it is specially designed for HDR evaluation by assessing structural similarity between fusion output and ground truth. Since MEF methods are mainly designed for static scenes, they lack the capability to handle motion, resulting in low TMQI scores and overlay artifacts, as shown in Fig. 9. HDR reconstruction methods trained on dynamic datasets achieve better performance, but still produce noticeable artifact. On the other side, due to our soft inpainting guidance, UltraFusion is much more robust to misalignment and occlusion and the fusion output contains almost no artifacts. It achieves the highest TMQI in Tab. 2 and the best visual result in Fig. 9.
Evaluation on our UltraFusion benchmark. At last, the evaluation on our benchmark validates the robustness of UltraFusion in the most challenging scenes. On all four non-reference metrics, our method outperforms compititors by a large margin, as shown in Tab. 2. The qualitative comparison in Fig. 8 is also consistent with the quantitative metrics.
For example, in the red zoom-in patch, integrating the highly bright sun from the under-exposed image into the over-exposed image is extremely challenging. Other methods fail to maintain the shape of the sun or preserve high contrast in the fused region, while our method naturally reconstructs the sun, preserving its appearance and visual-pleasing tone of the whole image. Moreover, we conduct a user study on our proposed benchmark. Specifically, we randomly select 20 scenes from our benchmark and invite 136 different users to participate. For each scene, each user is asked to compare our method with a randomly chosen baseline. The user study in Fig. 5 (b) indicates our method is more favored by users than competitors. This outcome aligns with the non-reference metric evaluation, showing that our method produces more natural images.
4.3 Ablation studies
Table 3: Ablation studies of three key components of our proposed UltraFusion on RealHDRV dataset [40].
[…]
Figure 10: Visual results of ablating each key component of our method. Each key component contributes to the final results.
To validate the effectiveness of our UltraFusion, we conduct ablation studies on key components, followed by an in-depth exploration of the designs of decompose-and-fuse control branch.
Key components. First, we perform the ablation study of the proposed key components, including the alignment strategy and decompose-and-fuse control branch, on the RealHDRV [40] dataset, as shown in Tab. 3. We first remove our alignment strategy, opting instead to train the model directly on the original dataset and input differently exposed image pairs without coarse alignment. Without the pre-alignment stage and the data synthesis pipeline designed for it, performance drops significantly in terms of TMQI. Fig. 10 (a) also shows that the model fails to implicitly align large motions. Then, we replace the decompose-and-fuse control branch (DFCB) with the vanilla ControlNet [64], and it fails to fuse features with large exposure differences, which leads to the loss of details (see  Fig. 10 (b)) and a decrease in TMQI.
Figure 11: Detailed ablation study on the design choices of decompose-and-fuse control branch.
Decompose-and-fuse control branch. We also evaluate how the form of soft guidance provided by the under-exposed image impacts the recovery of the highlight regions. When using the RGB under-exposed image $I_{ue}$ as guidance, the model will ignore some details in the output (Fig. 11 (c)). Replacing $I_{ue}$ with its structure information $S_{ue}$ retains more details (Fig. 11 (d)), but without generating vivid color. By incorporating both under-exposed color information $C_{ue}$ and structure information $S_{ue}$, the reconstructed results can maintain more details and color consistency (see Fig. 11 (e)). At last, multi-scale cross-attention further improves the fusion result (see Fig. 11 (f)).
4.4 Application on general image fusion
Figure 12: Extension to general fusion. Given totally different underexposed images as guidance (upper right corner), our method can also generate different fusion results.
One advantage of the UltraFusion is that it can be extended to a general image fusion, thanks to the flexibility of the proposed guided inpainting. To illustrate this potential, we explore one additional interesting demo to fuse two irrelevant images captured by different cameras at different locations. As shown in Fig. 12, UltraFusion successfully copies the moon (b) or blue sky (c) to the over exposed image, unlocking many interesting potential applications.
5 Conclusion
In this work, we introduce a novel approach to HDR imaging, tackling challenges presented by significant exposure differences and large motion. By modeling the fusion process as a guided inpainting problem and utilizing the under-exposed image for soft guidance, our method is robust with alignment errors and circumvents tone mapping, resulting in natural, artifact-free outputs. We also propose decompose-and-fuse control branch to improve feature modulation of ControlNet. Extensive experiments on existing datasets and our captured benchmark demonstrate the robustness and effectiveness of our method compared to previous HDR methods.
With extreme exposure differences and challenging non-rigid motion, occlusion mask estimation may introduce errors, causing the restoration of certain highlight regions degrades to single image HDR. While our method is able to use diffusion priors to restore these highlights, restoration without under-exposed information remains unreliable. A more exposure-robust optical flow algorithm is highly desirable, and we leave it for future work.